# open-sci-ref-0.01: open and reproducible reference baselines for language model and dataset comparison

### v2

@@ Appendix A Additional experimental results
A.1 Results per tasks on 1T tokens
Fig.˜4 shows the performance while training the 1.7B model for the top 3 datasets while increasing the budget to 1T tokens. The ranking remains consistent, Nemotron staying mostly on the top, followed or matched by DCLM and FineWeb-Edu (with the exception of Lambada, where DCLM dominates), while performance is going up as expected when increasing token scale.
Aligning models on the common compute axis for evaluation as done in Fig. 3 and Tab. 3 can provide a good overview where various learning procedures stand relative to each other, which is often difficult if not making compute that went into the training explicit. Proper scaling trend estimation requires the models across scales to be close to compute optimal. In this work, measured reference baseline points are not guaranteed to be close to compute optimal. Measurements that are executed for scaling law derivation can identify models trained close to compute optimal Pareto front. Such measurements give much more accurate predictions about learning procedure behavior than scaling trends presented here. This however require more dense measurements and hyperparameter tuning effort [34].
A.2 Further insights from multilingual benchmarks
We perform comparisons on two parallel multilingual benchmarks: MMMLU, which contains translated MMLU instructions by OpenAI and Belebele [38].
In Fig.˜6, we show the performance on Belebele while training a 1.7B model on Nemotron-CC-HQ, which echoes the surprising results seen on MMMLU. While the dataset is supposed to be English only, the model trained exhibits above random-chance performance (25%) on languages such as French, Spanish, Danish, or Portuguese. This is a surprising result which ought to be investigated. We hypothesize that it could come from the fact that the tolerance threshold used when filtering for English text lets further substantial data in another languages to slip into Nemontron dataset [10].
A.3 Further reference baselines
Learning rate schedules. To make sure that reference training shows similar performance when executed with various schedules, we check training with a cosine schedule to match close constant learning rate and cooldown schedule (WSD). We execute baseline training on the C4 dataset with 50B tokens, using both the cosine and WSD schedules, tuning hyperparameters for each schedule. We see matching performance for cosine and WSD on downstream evaluations.
@@ A.3 Further reference baselines
Appendix B Hyperparameter tuning
We perform hyperparameter tuning, measuring test loss on C4 and FineWeb-Edu, selecting hyperparameters that result in the lowest test loss on the datasets using 50B training runs with 1.3B model scale. We first experimented with a standard cosine learning rate schedule (Suppl. Tab. 5, tuning hyperparameters to match the HuggingFace reference models trained for 350B on C4 and FineWeb-Edu.
After achieving the match and thus checking the training procedure sanity, we switched to a constant learning + cooldown (WSD) schedule and tuned hyperparameters until we matched or outperformed results obtained with a cosine learning rate schedule on both 50B and 300B token budgets for a 1.3B model scale, which is still small enough for the tuning experiments to be affordable (Suppl. Tab. 4). The tuned WSD schedule training procedure was then used for all follow-up experiments and all model scales (0.13B, 0.4B, 1.3B, 1.7B). Tuned procedure uses weight decay of 0.05, global batch size of 4M tokens, cooldown of 20% of total training duration. Warmup and learning rate depend on the chosen total token budget. For larger token budgets >> 300B, a large warmup of 25 000 iterations turned out to give performance gains.
Tokens Global bs (tokens) Iters lr Warmup Cooldown (20%)
50B 2.65M 18839 $2\cdot 10^{-3}$ 6000 3767
50B 4.12M 11921 $4\cdot 10^{-3}$ 1000 2384
300B 2.09M 143052 $1\cdot 10^{-3}$ 5000 28610
300B 4.12M 72661 $4\cdot 10^{-3}$ 25000 14532
1T 4.12M 242204 $4\cdot 10^{-3}$ 25000 48440
Tokens Global bs (tokens) iters lr warmup
50B 2.65M 18839 3e-3 6000
50B 4.03M 12406 6e-3 1000
300B 2.09M 143052 3e-3 5000
Table 5: Hyperparameters for cosine lr schedule
Appendix C Distributed training on supercomputers
@@
Model (B) GPUs micro bs context length global bs (smpl/token) TFLOPS/ GPU/s Tokens/ GPU/s Run Time, A100 h (50B/300B/1T) 0.13 84 12 4096 1008/4.12M 114 87710 1.89/11.31/37.70 0.4 100 10 4096 1000/4.09M 157 44550 3.12/18.71/62.35 1.3 252 4 4096 1008/4.12M 172 16800 3.28/19.68/65.60 1.7 216 6 2048 1296/2.65M 167 18490 3.48/20.87/69.56 1.7 252 4 4096 1008/4.12M 184 14490 3.80/22.81/76.05 1.7 252 2 8192 504/4.12M 186 12320 4.47/26.84/89.47 1.7 252 1 16384 252/4.12M 201 10080 5.47/32.80/109.32
@@
Appendix D Author contributions

### v3

@@ Appendix A Additional experimental results
A.1 Further reference baselines
@@
Appendix B Distributed training on supercomputers
@@
Model (B) GPUs micro bs context length global bs (smpl/token) TFLOPS/ GPU/s Tokens/ GPU/s Run Time, h (50B/300B/1T) GPU A100 h (50B/300B/1T) 0.13 84 12 4096 1008/4.12M 114 87710 1.89/11.31/37.70 158/950/3167 0.4 100 10 4096 1000/4.09M 157 44550 3.12/18.71/62.35 312/1871/6235 1.3 252 4 4096 1008/4.12M 172 16800 3.28/19.68/65.60 827/4959/16530 1.7 216 6 2048 1296/2.65M 167 18490 3.48/20.87/69.56 751/4508/15026 1.7 252 4 4096 1008/4.12M 184 14490 3.80/22.81/76.05 958/5749/19164 1.7 252 2 8192 504/4.12M 186 12320 4.47/26.84/89.47 1127/6764/22546 1.7 252 1 16384 252/4.12M 201 10080 5.47/32.80/109.32 1377/8265/27549
@@
Appendix C Author contributions